Concord Company uses the gross profit method to estimate inventory for monthly reporting purposes. Presented below is information for the month of May.

Compute the estimated inventory at May 31, assuming that the gross profit is 25% of cost. (Round percentage of sales to 2 decimal places, e.g. 78.74% and final answer to 0 decimal places, e.g. 6,225.) ________

Inventory, May 1 $ 149,100
Purchases (gross) 683,200
Freight-in 28,700
Sales revenue 998,600
Sales returns 76,500
Purchase discounts 11,000
Answer #1

Cost of goods sold = net sales - gross profit

Cost of goods sold = (998600-76500) - 25% cost of goods sold

Cost of goods sold = 922100/1.25

Cost of goods sold = $737680

Ending inventory = beginning inventory + purchases net + freight - cost of goods sold

= 149100 + (683200-11000)+28700 - 737680

= $112320

Thus the correct answer is $ 112320